OpenStudy (anonymous):

Simplify using order of operations: 2 x (3 – 6) + 12. A. 6 B. 9 C. 12 D. 18

OpenStudy (calculusxy):

Do you know the steps of the Order of Operations?

OpenStudy (anonymous):

\[PEDMAS\]

OpenStudy (calculusxy):

Well I believe that it is PEMDAS. You have to do Multiplication and Division left to right and Addition and Subtraction also .

OpenStudy (anonymous):

Still don't understand it

jigglypuff314 (jigglypuff314):

PEMDAS Parenthesis Exponents Multiplication/Division (left to right) Addition/Subtraction (left to right)

OpenStudy (calculusxy):

P for parenthesis E for exponents M for multiplication D for division A for addition S for subtraction *Multiplication and Division is done left to right. Meaning whatever is on the left do that one first. *Addition and Subtraction is done left to right.

jigglypuff314 (jigglypuff314):

so you would do the stuff in the parenthesis first ( stuff ) so for 2 x (3 – 6) + 12 you would first do 3 - 6 = -( 6 - 3) = ?

OpenStudy (calculusxy):

Follow the orders of the operation and you will arrive with your answer. Don't worry, I will help you on answering this question.

OpenStudy (calculusxy):

@musa19965 What do you do first based on the PEMDAS?

jigglypuff314 (jigglypuff314):

@xionfairy first of all, you're not supposed to give answers second of all, you forgot a negative and so your answer is wrong...

OpenStudy (calculusxy):

\[2\times(3-6)+12\] Based on this, do you see a parenthesis (since parenthesis is first in the Order of Operations)?

OpenStudy (anonymous):

@jigglypuff314 apologies!! i really wasn't paying attention!

OpenStudy (calculusxy):

@xionfairy Don't worry everyone makes mistakes :)

OpenStudy (anonymous):

man idk this pellet

jigglypuff314 (jigglypuff314):

so you would do the stuff in the parenthesis first ( stuff ) so for 2 x (3 – 6) + 12 you would first do 3 - 6 = -( 6 - 3) = ?

OpenStudy (calculusxy):

( ) These two symbols are parenthesis. Do you see this in any part of the equation?

OpenStudy (shamil98):

2 x (3-6) + 12 you do the subtraction the parantheses first (3-6) you then are left with 2 x (-3) + 12 then you do your multiplication. 2 x - 3 = -6 finally you are left with addition -6 + 12, keep in mind this is the same as 12 - 6 your answer is 6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

 @jigglypuff314 better?